a and b total 32 between them
b and c total 33 between them
c and d total 28 between them
d and e total 7 between them
a,b,c,d,e total between them 65
what is a,b,c,d,e

Answer by solver91311(24713)



a + b             = 32
    b + c         = 33
        c + d     = 28
            d + e =  7
a + b + c + d + e = 65

Solve the 5X5 system for and . Hint: Substitute the value of and into the fifth equation and solve for the value of , from which you can get , from which you can get ...

John

My calculator said it, I believe it, that settles it